OpenStudy (anonymous):

use the half-angle formula to find the exact value of sin5*pi/12.. help?

OpenStudy (freckles):

hint \[\sin(\frac{5\pi}{6} \cdot \frac{1}{2})\]

OpenStudy (freckles):

\[\sin^2(\frac{1}{2}x)=\frac{1-\cos(x)}{2}\]